\begin{thebibliography}{99}

%\cite{Batley:2002gn}
\bibitem{Batley:2002gn}
  J.~R.~Batley {\it et al.} [NA48 Collaboration],
  %``A Precision measurement of direct CP violation in the decay of neutral kaons into two pions,''
  Phys.\ Lett.\ B {\bf 544} (2002) 97
  %doi:10.1016/S0370-2693(02)02476-0
  [hep-ex/0208009].
  %%CITATION = doi:10.1016/S0370-2693(02)02476-0;%%
  %353 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{AlaviHarati:2002ye}
\bibitem{AlaviHarati:2002ye}
  A.~Alavi-Harati {\it et al.} [KTeV Collaboration],
  %``Measurements of direct CP violation, CPT symmetry, and other parameters in the neutral kaon system,''
  Phys.\ Rev.\ D {\bf 67} (2003) 012005
   Erratum: [Phys.\ Rev.\ D {\bf 70} (2004) 079904]
  %doi:10.1103/PhysRevD.70.079904, 10.1103/PhysRevD.67.012005
  [hep-ex/0208007].
  %%CITATION = doi:10.1103/PhysRevD.70.079904, 10.1103/PhysRevD.67.012005;%%
  %368 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Abouzaid:2010ny}
\bibitem{Abouzaid:2010ny}
  E.~Abouzaid {\it et al.} [KTeV Collaboration],
  %``Precise Measurements of Direct CP Violation, CPT Symmetry, and Other Parameters in the Neutral Kaon System,''
  Phys.\ Rev.\ D {\bf 83} (2011) 092001
  %doi:10.1103/PhysRevD.83.092001
  [arXiv:1011.0127 [hep-ex]].
  %%CITATION = doi:10.1103/PhysRevD.83.092001;%%
  %135 citations counted in INSPIRE as of 05 Dec 2020
  
   %\cite{Buras:1993dy}
\bibitem{Buras:1993dy}
  A.~J.~Buras, M.~Jamin and M.~E.~Lautenbacher,
  %``The Anatomy of $\epsilon' / \epsilon$ beyond leading logarithms with improved hadronic matrix elements,''
  Nucl.\ Phys.\ B {\bf 408} (1993) 209
  %doi:10.1016/0550-3213(93)90535-W
  [hep-ph/9303284].
  %%CITATION = doi:10.1016/0550-3213(93)90535-W;%%
  %382 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Buras:1996dq}
\bibitem{Buras:1996dq}
  A.~J.~Buras, M.~Jamin and M.~E.~Lautenbacher,
  %``A 1996 analysis of the CP violating ratio epsilon-prime / epsilon,''
  Phys.\ Lett.\ B {\bf 389} (1996) 749
  %doi:10.1016/S0370-2693(96)80019-0, 10.1016/S0370-2693(96)01349-4
  [hep-ph/9608365].
  %%CITATION = doi:10.1016/S0370-2693(96)80019-0, 10.1016/S0370-2693(96)01349-4;%%
  %133 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Bosch:1999wr}
\bibitem{Bosch:1999wr}
  S.~Bosch, A.~J.~Buras, M.~Gorbahn, S.~Jager, M.~Jamin, M.~E.~Lautenbacher and L.~Silvestrini,
  %``Standard model confronting new results for epsilon-prime / epsilon,''
  Nucl.\ Phys.\ B {\bf 565} (2000) 3
  %doi:10.1016/S0550-3213(99)00694-X
  [hep-ph/9904408].
  %%CITATION = doi:10.1016/S0550-3213(99)00694-X;%%
  %179 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Buras:2000qz}
\bibitem{Buras:2000qz}
  A.~J.~Buras, P.~Gambino, M.~Gorbahn, S.~Jager and L.~Silvestrini,
  %``epsilon-prime / epsilon and rare K and B decays in the MSSM,''
  Nucl.\ Phys.\ B {\bf 592} (2001) 55
  %doi:10.1016/S0550-3213(00)00582-4
  [hep-ph/0007313].
  %%CITATION = doi:10.1016/S0550-3213(00)00582-4;%%
  %174 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Ciuchini:1995cd}
\bibitem{Ciuchini:1995cd}
  M.~Ciuchini, E.~Franco, G.~Martinelli, L.~Reina and L.~Silvestrini,
  %``An Upgraded analysis of epsilon-prime epsilon at the next-to-leading order,''
  Z.\ Phys.\ C {\bf 68} (1995) 239
  %doi:10.1007/BF01566672
  [hep-ph/9501265].
  %%CITATION = doi:10.1007/BF01566672;%%
  %223 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Ciuchini:1992tj}
\bibitem{Ciuchini:1992tj}
  M.~Ciuchini, E.~Franco, G.~Martinelli and L.~Reina,
  %``$\epsilon' / \epsilon$ at the Next-to-leading order in QCD and QED,''
  Phys.\ Lett.\ B {\bf 301} (1993) 263
  %doi:10.1016/0370-2693(93)90699-I
  [hep-ph/9212203].
  %%CITATION = doi:10.1016/0370-2693(93)90699-I;%%
  %259 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Pallante:1999qf}
\bibitem{Pallante:1999qf}
  E.~Pallante and A.~Pich,
  %``Strong enhancement of epsilon-prime / epsilon through final state interactions,''
  Phys.\ Rev.\ Lett.\  {\bf 84} (2000) 2568
  %doi:10.1103/PhysRevLett.84.2568
  [hep-ph/9911233].
  %%CITATION = doi:10.1103/PhysRevLett.84.2568;%%
  %161 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Pallante:2000hk}
\bibitem{Pallante:2000hk}
  E.~Pallante and A.~Pich,
  %``Final state interactions in kaon decays,''
  Nucl.\ Phys.\ B {\bf 592} (2001) 294
  %doi:10.1016/S0550-3213(00)00601-5
  [hep-ph/0007208].
  %%CITATION = doi:10.1016/S0550-3213(00)00601-5;%%
  %126 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Pallante:2001he}
\bibitem{Pallante:2001he}
  E.~Pallante, A.~Pich and I.~Scimemi,
  %``The Standard model prediction for epsilon-prime / epsilon,''
  Nucl.\ Phys.\ B {\bf 617} (2001) 441
  %doi:10.1016/S0550-3213(01)00418-7
  [hep-ph/0105011].
  %%CITATION = doi:10.1016/S0550-3213(01)00418-7;%%
  %136 citations counted in INSPIRE as of 05 Dec 2020
  
    %\cite{Blum:2011ng}
\bibitem{Blum:2011ng}
  T.~Blum {\it et al.},
  %``The $K\to(\pi\pi)_{I=2}$ Decay Amplitude from Lattice QCD,''
  Phys.\ Rev.\ Lett.\  {\bf 108} (2012) 141601
  %doi:10.1103/PhysRevLett.108.141601
  [arXiv:1111.1699 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevLett.108.141601;%%
  %123 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Blum:2012uk}
\bibitem{Blum:2012uk}
  T.~Blum {\it et al.},
  %``Lattice determination of the $K \to (\pi\pi)_{I=2}$ Decay Amplitude $A_2$,''
  Phys.\ Rev.\ D {\bf 86} (2012) 074513
  %doi:10.1103/PhysRevD.86.074513
  [arXiv:1206.5142 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.86.074513;%%
  %117 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Blum:2015ywa}
\bibitem{Blum:2015ywa}
  T.~Blum {\it et al.},
  %``$K \rightarrow \pi\pi$ $\Delta I=3/2$ decay amplitude in the continuum limit,''
  Phys.\ Rev.\ D {\bf 91} (2015) no.7,  074502
  %doi:10.1103/PhysRevD.91.074502
  [arXiv:1502.00263 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.91.074502;%%
  %143 citations counted in INSPIRE as of 05 Dec 2020
  
    %\cite{Boyle:2012ys}
\bibitem{Boyle:2012ys}
  P.~A.~Boyle {\it et al.} [RBC and UKQCD Collaborations],
  %``Emerging understanding of the $\Delta I = 1/2$ Rule from Lattice QCD,''
  Phys.\ Rev.\ Lett.\  {\bf 110} (2013) no.15,  152001
  %doi:10.1103/PhysRevLett.110.152001
  [arXiv:1212.1474 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevLett.110.152001;%%
  %91 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Pich:1995qp}
\bibitem{Pich:1995qp}
  A.~Pich and E.~de Rafael,
  %``Weak K amplitudes in the chiral and 1/n(c) expansions,''
  Phys.\ Lett.\ B {\bf 374} (1996) 186
  %doi:10.1016/0370-2693(96)00171-2
  [hep-ph/9511465].
  %%CITATION = doi:10.1016/0370-2693(96)00171-2;%%
  %57 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Pich:1990mw}
\bibitem{Pich:1990mw}
  A.~Pich and E.~de Rafael,
  %``Four quark operators and nonleptonic weak transitions,''
  Nucl.\ Phys.\ B {\bf 358} (1991) 311.
  %doi:10.1016/0550-3213(91)90351-W
  %%CITATION = doi:10.1016/0550-3213(91)90351-W;%%
  %207 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Jamin:1994sv}
\bibitem{Jamin:1994sv}
  M.~Jamin and A.~Pich,
  %``QCD corrections to inclusive Delta S = 1,2 transitions at the next-to-leading order,''
  Nucl.\ Phys.\ B {\bf 425} (1994) 15
  %doi:10.1016/0550-3213(94)90171-6
  [hep-ph/9402363].
  %%CITATION = doi:10.1016/0550-3213(94)90171-6;%%
  %48 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Bertolini:1997ir}
\bibitem{Bertolini:1997ir}
  S.~Bertolini, J.~O.~Eeg, M.~Fabbrichesi and E.~I.~Lashin,
  %``The Delta I = 1/2 rule and B(K) at O (p**4) in the chiral expansion,''
  Nucl.\ Phys.\ B {\bf 514} (1998) 63
  %doi:10.1016/S0550-3213(97)00787-6
  [hep-ph/9705244].
  %%CITATION = doi:10.1016/S0550-3213(97)00787-6;%%
  %103 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Antonelli:1995nv}
\bibitem{Antonelli:1995nv}
  V.~Antonelli, S.~Bertolini, J.~O.~Eeg, M.~Fabbrichesi and E.~I.~Lashin,
  %``The Delta S = 1 weak chiral lagrangian as the effective theory of the chiral quark model,''
  Nucl.\ Phys.\ B {\bf 469} (1996) 143
  %doi:10.1016/0550-3213(96)00144-7
  [hep-ph/9511255].
  %%CITATION = doi:10.1016/0550-3213(96)00144-7;%%
  %64 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Antonelli:1995gw}
\bibitem{Antonelli:1995gw}
  V.~Antonelli, S.~Bertolini, M.~Fabbrichesi and E.~I.~Lashin,
  %``The Delta I = 1/2 selection rule,''
  Nucl.\ Phys.\ B {\bf 469} (1996) 181
  %doi:10.1016/0550-3213(96)00145-9
  [hep-ph/9511341].
  %%CITATION = doi:10.1016/0550-3213(96)00145-9;%%
  %47 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Hambye:1998sma}
\bibitem{Hambye:1998sma}
  T.~Hambye, G.~O.~Kohler, E.~A.~Paschos, P.~H.~Soldan and W.~A.~Bardeen,
  %``1 / N(c) corrections to the hadronic matrix elements of Q(6) and Q(8) in K ---> pi pi decays,''
  Phys.\ Rev.\ D {\bf 58} (1998) 014017
  %doi:10.1103/PhysRevD.58.014017
  [hep-ph/9802300].
  %%CITATION = doi:10.1103/PhysRevD.58.014017;%%
  %122 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Bardeen:1986vz}
\bibitem{Bardeen:1986vz}
  W.~A.~Bardeen, A.~J.~Buras and J.~M.~Gerard,
  %``A Consistent Analysis of the Delta I = 1/2 Rule for K Decays,''
  Phys.\ Lett.\ B {\bf 192} (1987) 138.
  %doi:10.1016/0370-2693(87)91156-7
  %%CITATION = doi:10.1016/0370-2693(87)91156-7;%%
  %350 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Buras:2014maa}
\bibitem{Buras:2014maa}
  A.~J.~Buras, J.~M.~Gérard and W.~A.~Bardeen,
  %``Large $N$ Approach to Kaon Decays and Mixing 28 Years Later: $\Delta I = 1/2$ Rule, $\hat B_K$ and $\Delta M_K$,''
  Eur.\ Phys.\ J.\ C {\bf 74} (2014) 2871
  %doi:10.1140/epjc/s10052-014-2871-x
  [arXiv:1401.1385 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-014-2871-x;%%
  %97 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Bijnens:1998ee}
\bibitem{Bijnens:1998ee}
  J.~Bijnens and J.~Prades,
  %``The Delta I = 1/2 rule in the chiral limit,''
  JHEP {\bf 9901} (1999) 023
  %doi:10.1088/1126-6708/1999/01/023
  [hep-ph/9811472].
  %%CITATION = doi:10.1088/1126-6708/1999/01/023;%%
  %93 citations counted in INSPIRE as of 05 Dec 2020
  
   %\cite{Bai:2015nea}
\bibitem{Bai:2015nea}
  Z.~Bai {\it et al.} [RBC and UKQCD Collaborations],
  %``Standard Model Prediction for Direct CP Violation in K→ππ Decay,''
  Phys.\ Rev.\ Lett.\  {\bf 115} (2015) no.21,  212001
  %doi:10.1103/PhysRevLett.115.212001
  [arXiv:1505.07863 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevLett.115.212001;%%
  %197 citations counted in INSPIRE as of 05 Dec 2020
  
    %\cite{Gisbert:2018tuf}
\bibitem{Gisbert:2018tuf}
  H.~Gisbert and A.~Pich,
  %``Updated Standard Model Prediction for $\varepsilon'/\varepsilon$,''
  Nucl.\ Part.\ Phys.\ Proc.\  {\bf 300-302} (2018) 137
  %doi:10.1016/j.nuclphysbps.2018.12.024
  [arXiv:1810.04904 [hep-ph]].
  %%CITATION = doi:10.1016/j.nuclphysbps.2018.12.024;%%
  %8 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Gisbert:2018niu}
\bibitem{Gisbert:2018niu}
  H.~Gisbert,
  %``Current status of $\varepsilon'/\varepsilon$ in the Standard Model,''
  PoS Confinement {\bf 2018} (2018) 178
  %doi:10.22323/1.336.0178
  [arXiv:1811.12206 [hep-ph]].
  %%CITATION = doi:10.22323/1.336.0178;%%
  %3 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Gisbert:2019uyx}
\bibitem{Gisbert:2019uyx}
  H.~Gisbert,
  ``Status of $\epsilon'$ /$\epsilon$ in the Standard Model'', Moriond EW 2019, 277-282.
  %%CITATION = INSPIRE-1800634;%%
  
  %\cite{Buras:2018ozh}
\bibitem{Buras:2018ozh}
  A.~J.~Buras,
  %``$\epsilon^\prime/\epsilon$-2018: A Christmas Story,''
  arXiv:1812.06102 [hep-ph].
  %%CITATION = ARXIV:1812.06102;%%
  %9 citations counted in INSPIRE as of 05 Dec 2020
  
    %\cite{Buras:2015xba}
\bibitem{Buras:2015xba}
  A.~J.~Buras and J.~M.~Gérard,
  %``Upper bounds on ε′/ε parameters B$_{6}^{(1/2)}$ and B$_{8}^{(3/2)}$ from large N QCD and other news,''
  JHEP {\bf 1512} (2015) 008
  %doi:10.1007/JHEP12(2015)008
  [arXiv:1507.06326 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP12(2015)008;%%
  %81 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Buras:2016fys}
\bibitem{Buras:2016fys}
  A.~J.~Buras and J.~M.~Gerard,
  %``Final state interactions in $K\rightarrow \pi \pi $ decays: $\Delta I=1/2$ rule vs. $\varepsilon '/\varepsilon $,''
  Eur.\ Phys.\ J.\ C {\bf 77} (2017) no.1,  10
  %doi:10.1140/epjc/s10052-016-4586-7
  [arXiv:1603.05686 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-016-4586-7;%%
  %65 citations counted in INSPIRE as of 05 Dec 2020
  
  %\cite{Buras:2015yba}
\bibitem{Buras:2015yba}
  A.~J.~Buras, M.~Gorbahn, S.~Jäger and M.~Jamin,
  %``Improved anatomy of ε′/ε in the Standard Model,''
  JHEP {\bf 1511} (2015) 202
  %doi:10.1007/JHEP11(2015)202
  [arXiv:1507.06345 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP11(2015)202;%%
  %125 citations counted in INSPIRE as of 05 Dec 2020
  
    %\cite{Gisbert:2017vvj}
\bibitem{Gisbert:2017vvj}
  H.~Gisbert and A.~Pich,
  %``Direct CP violation in $K^0\to\pi\pi$: Standard Model Status,''
  Rept.\ Prog.\ Phys.\  {\bf 81} (2018) no.7,  076201
  %doi:10.1088/1361-6633/aac18e
  [arXiv:1712.06147 [hep-ph]].
  %%CITATION = doi:10.1088/1361-6633/aac18e;%%
  %49 citations counted in INSPIRE as of 05 Dec 2020
  
    %\cite{Cirigliano:2019cpi}
\bibitem{Cirigliano:2019cpi}
  V.~Cirigliano, H.~Gisbert, A.~Pich and A.~Rodríguez-Sánchez,
  %``Isospin-violating contributions to  $\epsilon'/\epsilon$,''
  JHEP {\bf 2002} (2020) 032
  %doi:10.1007/JHEP02(2020)032
  [arXiv:1911.01359 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP02(2020)032;%%
  %14 citations counted in INSPIRE as of 06 Dec 2020
  
   %\cite{Cirigliano:2019obu}
\bibitem{Cirigliano:2019obu}
  V.~Cirigliano, H.~Gisbert, A.~Pich and A.~Rodríguez-Sánchez,
  %``A complete update of $\varepsilon'/\varepsilon$ in the Standard Model,''
  PoS EPS {\bf -HEP2019} (2020) 238
  %doi:10.22323/1.364.0238
  [arXiv:1911.06554 [hep-ph]].
  %%CITATION = doi:10.22323/1.364.0238;%%
  %2 citations counted in INSPIRE as of 06 Dec 2020
  
  %\cite{Cirigliano:2019zjv}
\bibitem{Cirigliano:2019zjv}
  V.~Cirigliano, H.~Gisbert, A.~Pich and A.~Rodríguez-Sánchez,
  %``Isospin-breaking contributions to $\varepsilon'/\varepsilon$,''
  J.\ Phys.\ Conf.\ Ser.\  {\bf 1526} (2020) 012010
  %doi:10.1088/1742-6596/1526/1/012010
  [arXiv:1912.04811 [hep-ph]].
  %%CITATION = doi:10.1088/1742-6596/1526/1/012010;%%
  %1 citations counted in INSPIRE as of 06 Dec 2020
  
  %\cite{Cirigliano:2019ani}
\bibitem{Cirigliano:2019ani}
  V.~Cirigliano, H.~Gisbert, A.~Pich and A.~Rodríguez-Sánchez,
  %``Theoretical status of $\varepsilon'/\varepsilon$,''
  J.\ Phys.\ Conf.\ Ser.\  {\bf 1526} (2020) 012011
  %doi:10.1088/1742-6596/1526/1/012011
  [arXiv:1912.04736 [hep-ph]].
  %%CITATION = doi:10.1088/1742-6596/1526/1/012011;%%
  %2 citations counted in INSPIRE as of 06 Dec 2020
  
    %\cite{Abbott:2020hxn}
\bibitem{Abbott:2020hxn}
  R.~Abbott {\it et al.} [RBC and UKQCD Collaborations],
  %``Direct CP violation and the $\Delta I=1/2$ rule in $K\to\pi\pi$ decay from the standard model,''
  Phys.\ Rev.\ D {\bf 102} (2020) no.5,  054509
  %doi:10.1103/PhysRevD.102.054509
  [arXiv:2004.09440 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.102.054509;%%
  %9 citations counted in INSPIRE as of 06 Dec 2020
  
   %\cite{Cirigliano:2003nn}
\bibitem{Cirigliano:2003nn}
  V.~Cirigliano, A.~Pich, G.~Ecker and H.~Neufeld,
  %``Isospin violation in epsilon-prime,''
  Phys.\ Rev.\ Lett.\  {\bf 91} (2003) 162001
  %doi:10.1103/PhysRevLett.91.162001
  [hep-ph/0307030].
  %%CITATION = doi:10.1103/PhysRevLett.91.162001;%%
  %78 citations counted in INSPIRE as of 06 Dec 2020
  
  %\cite{Cirigliano:2003gt}
\bibitem{Cirigliano:2003gt}
  V.~Cirigliano, G.~Ecker, H.~Neufeld and A.~Pich,
  %``Isospin breaking in K ---> pi pi decays,''
  Eur.\ Phys.\ J.\ C {\bf 33} (2004) 369
  %doi:10.1140/epjc/s2003-01579-3
  [hep-ph/0310351].
  %%CITATION = doi:10.1140/epjc/s2003-01579-3;%%
  %91 citations counted in INSPIRE as of 06 Dec 2020
  
    %\cite{Aebischer:2020jto}
\bibitem{Aebischer:2020jto}
  J.~Aebischer, C.~Bobeth and A.~J.~Buras,
  %``$\varepsilon '/\varepsilon $ in the Standard Model at the Dawn of the 2020s,''
  Eur.\ Phys.\ J.\ C {\bf 80} (2020) no.8,  705
  %doi:10.1140/epjc/s10052-020-8267-1
  [arXiv:2005.05978 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-020-8267-1;%%
  %4 citations counted in INSPIRE as of 06 Dec 2020
  
    %\cite{Antonelli:2010yf}
\bibitem{Antonelli:2010yf}
  M.~Antonelli {\it et al.} [FlaviaNet Working Group on Kaon Decays],
  %``An Evaluation of $|V_{us}|$ and precise tests of the Standard Model from world data on leptonic and semileptonic kaon decays,''
  Eur.\ Phys.\ J.\ C {\bf 69} (2010) 399
  %doi:10.1140/epjc/s10052-010-1406-3
  [arXiv:1005.2323 [hep-ph]].
  %%CITATION = doi:10.1140/epjc/s10052-010-1406-3;%%
  %290 citations counted in INSPIRE as of 06 Dec 2020
  
  %\cite{Buchalla:1995vs}
\bibitem{Buchalla:1995vs}
  G.~Buchalla, A.~J.~Buras and M.~E.~Lautenbacher,
  %``Weak decays beyond leading logarithms,''
  Rev.\ Mod.\ Phys.\  {\bf 68} (1996) 1125
  %doi:10.1103/RevModPhys.68.1125
  [hep-ph/9512380].
  %%CITATION = doi:10.1103/RevModPhys.68.1125;%%
  %2599 citations counted in INSPIRE as of 09 Dec 2020
  
  %\cite{Buras:1991jm}
\bibitem{Buras:1991jm}
  A.~J.~Buras, M.~Jamin, M.~E.~Lautenbacher and P.~H.~Weisz,
  %``Effective Hamiltonians for $\Delta S = 1$ and $\Delta B = 1$ nonleptonic decays beyond the leading logarithmic approximation,''
  Nucl.\ Phys.\ B {\bf 370} (1992) 69
   Addendum: [Nucl.\ Phys.\ B {\bf 375} (1992) 501].
  %doi:10.1016/0550-3213(92)90345-C
  %%CITATION = doi:10.1016/0550-3213(92)90345-C;%%
  %450 citations counted in INSPIRE as of 09 Dec 2020
  
%\cite{Buras:1992tc}
\bibitem{Buras:1992tc}
  A.~J.~Buras, M.~Jamin, M.~E.~Lautenbacher and P.~H.~Weisz,
  %``Two loop anomalous dimension matrix for $\Delta S = 1$ weak nonleptonic decays I: $\mathcal{O}(\alpha_s^2)$,''
  Nucl.\ Phys.\ B {\bf 400} (1993) 37
  %doi:10.1016/0550-3213(93)90397-8
  [hep-ph/9211304].
  %%CITATION = doi:10.1016/0550-3213(93)90397-8;%%
  %354 citations counted in INSPIRE as of 09 Dec 2020
  
%\cite{Buras:1992zv}
\bibitem{Buras:1992zv}
  A.~J.~Buras, M.~Jamin and M.~E.~Lautenbacher,
  %``Two loop anomalous dimension matrix for Delta S = 1 weak nonleptonic decays. 2. O(alpha-alpha-s),''
  Nucl.\ Phys.\ B {\bf 400} (1993) 75
  %doi:10.1016/0550-3213(93)90398-9
  [hep-ph/9211321].
  %%CITATION = doi:10.1016/0550-3213(93)90398-9;%%
  %281 citations counted in INSPIRE as of 09 Dec 2020
  
%\cite{Ciuchini:1993vr}
\bibitem{Ciuchini:1993vr}
  M.~Ciuchini, E.~Franco, G.~Martinelli and L.~Reina,
  %``The Delta S = 1 effective Hamiltonian including next-to-leading order QCD and QED corrections,''
  Nucl.\ Phys.\ B {\bf 415} (1994) 403
  %doi:10.1016/0550-3213(94)90118-X
  [hep-ph/9304257].
  %%CITATION = doi:10.1016/0550-3213(94)90118-X;%%
  %500 citations counted in INSPIRE as of 09 Dec 2020
  
    %\cite{Buras:1999st}
\bibitem{Buras:1999st}
  A.~J.~Buras, P.~Gambino and U.~A.~Haisch,
  %``Electroweak penguin contributions to nonleptonic Delta F = 1 decays at NNLO,''
  Nucl.\ Phys.\ B {\bf 570} (2000) 117
  %doi:10.1016/S0550-3213(99)00810-X
  [hep-ph/9911250].
  %%CITATION = doi:10.1016/S0550-3213(99)00810-X;%%
  %71 citations counted in INSPIRE as of 09 Dec 2020
  
  %\cite{Gorbahn:2004my}
\bibitem{Gorbahn:2004my}
  M.~Gorbahn and U.~Haisch,
  %``Effective Hamiltonian for non-leptonic $|\Delta F| = 1$ decays at NNLO in QCD,''
  Nucl.\ Phys.\ B {\bf 713} (2005) 291
  %doi:10.1016/j.nuclphysb.2005.01.047
  [hep-ph/0411071].
  %%CITATION = doi:10.1016/j.nuclphysb.2005.01.047;%%
  %229 citations counted in INSPIRE as of 09 Dec 2020
  
  %\cite{Cerda-Sevilla:2016yzo}
\bibitem{Cerda-Sevilla:2016yzo}
  M.~Cerdà-Sevilla, M.~Gorbahn, S.~Jäger and A.~Kokulu,
  %``Towards NNLO accuracy for ε′/ε,''
  J.\ Phys.\ Conf.\ Ser.\  {\bf 800} (2017) no.1,  012008
  %doi:10.1088/1742-6596/800/1/012008
  [arXiv:1611.08276 [hep-ph]].
  %%CITATION = doi:10.1088/1742-6596/800/1/012008;%%
  %22 citations counted in INSPIRE as of 09 Dec 2020
  
  %\cite{Sanchez:2018atj}
\bibitem{Sanchez:2018atj}
  A.~Rodríguez-Sánchez and A.~Pich,
  %``Confronting hadronic tau decays with non-leptonic kaon decays,''
  Nucl.\ Part.\ Phys.\ Proc.\  {\bf 300-302} (2018) 131
  %doi:10.1016/j.nuclphysbps.2018.12.023
  [arXiv:1811.06706 [hep-ph]].
  %%CITATION = doi:10.1016/j.nuclphysbps.2018.12.023;%%
  %2 citations counted in INSPIRE as of 10 Dec 2020  
  
%\cite{AntonioS}
\bibitem{AntonioS}
    A.~Pich and A.~Rodríguez-Sánchez, in preparation.
    
    
    %\cite{Buchler:2001nm}
\bibitem{Buchler:2001nm}
  M.~Buchler, G.~Colangelo, J.~Kambor and F.~Orellana,
  %``Dispersion relations and soft pion theorems for K ---> pi pi,''
  Phys.\ Lett.\ B {\bf 521} (2001) 22
  %doi:10.1016/S0370-2693(01)01098-X
  [hep-ph/0102287].
  %%CITATION = doi:10.1016/S0370-2693(01)01098-X;%%
  %35 citations counted in INSPIRE as of 10 Dec 2020
  
  %\cite{Buchler:2005xn}
\bibitem{Buchler:2005xn}
  M.~Buchler,
  %``The Chiral logs of the K ---> pi pi amplitude,''
  Phys.\ Lett.\ B {\bf 633} (2006) 497
  %doi:10.1016/j.physletb.2005.12.018
  [hep-ph/0511087].
  %%CITATION = doi:10.1016/j.physletb.2005.12.018;%%
  %9 citations counted in INSPIRE as of 10 Dec 2020
  
    %\cite{Bertolini:1995tp}
\bibitem{Bertolini:1995tp}
  S.~Bertolini, J.~O.~Eeg and M.~Fabbrichesi,
  %``A New estimate of epsilon-prime / epsilon,''
  Nucl.\ Phys.\ B {\bf 476} (1996) 225
  %doi:10.1016/0550-3213(96)00335-5
  [hep-ph/9512356].
  %%CITATION = doi:10.1016/0550-3213(96)00335-5;%%
  %67 citations counted in INSPIRE as of 10 Dec 2020
  
%\cite{Bertolini:1997nf}
\bibitem{Bertolini:1997nf}
  S.~Bertolini, J.~O.~Eeg, M.~Fabbrichesi and E.~I.~Lashin,
  %``Epsilon-prime / epsilon at O(p**4) in the chiral expansion,''
  Nucl.\ Phys.\ B {\bf 514} (1998) 93
  %doi:10.1016/S0550-3213(97)00786-4
  [hep-ph/9706260].
  %%CITATION = doi:10.1016/S0550-3213(97)00786-4;%%
  %160 citations counted in INSPIRE as of 10 Dec 2020
  
%\cite{Bertolini:1998vd}
\bibitem{Bertolini:1998vd}
  S.~Bertolini, M.~Fabbrichesi and J.~O.~Eeg,
  %``Theory of the CP Violating Parameter $\epsilon^{\prime} / \epsilon$,''
  Rev.\ Mod.\ Phys.\  {\bf 72} (2000) 65
  %doi:10.1103/RevModPhys.72.65
  [hep-ph/9802405].
  %%CITATION = doi:10.1103/RevModPhys.72.65;%%
  %189 citations counted in INSPIRE as of 10 Dec 2020
  
  %\cite{Bijnens:2000im}
\bibitem{Bijnens:2000im}
  J.~Bijnens and J.~Prades,
  %``Epsilon-prime K / K epsilon in the chiral limit,''
  JHEP {\bf 0006} (2000) 035
  %doi:10.1088/1126-6708/2000/06/035
  [hep-ph/0005189].
  %%CITATION = doi:10.1088/1126-6708/2000/06/035;%%
  %108 citations counted in INSPIRE as of 10 Dec 2020

%\cite{Hambye:2003cy}
\bibitem{Hambye:2003cy}
  T.~Hambye, S.~Peris and E.~de Rafael,
  %``Delta I = 1/2 and epsilon-prime / epsilon in large N(c) QCD,''
  JHEP {\bf 0305} (2003) 027
  %doi:10.1088/1126-6708/2003/05/027
  [hep-ph/0305104].
  %%CITATION = doi:10.1088/1126-6708/2003/05/027;%%
  %83 citations counted in INSPIRE as of 10 Dec 2020
  
  
\end{thebibliography}

\documentclass[3p,times,twocolumn]{elsarticle}
\biboptions{comma,sort&compress}
\usepackage{graphicx}
\usepackage[utf8]{inputenc}
\usepackage{here}
\usepackage{color}
\usepackage{ecrc}

\usepackage{tikz}
\usetikzlibrary{positioning,arrows}
\usetikzlibrary{decorations.pathmorphing}
\usetikzlibrary{decorations.markings}
\usepackage{pgfplots}
\usepackage{xparse}
\usetikzlibrary{positioning,arrows,patterns}
\usetikzlibrary{decorations.markings}
\usetikzlibrary{calc}

\volume{00}

\firstpage{1}

\journalname{Nuclear and Particle Physics Proceedings}

\runauth{}

\jid{nppp}

\jnltitlelogo{Nuclear and Particle Physics Proceedings}

\usepackage{amssymb}

\usepackage[figuresright]{rotating}

\def\bc{\begin{center}}
\def\ec{\end{center}}
\def\be{\begin{equation}}
\def\ee{\end{equation}}
\def\beqn{\begin{eqnarray}}
\def\eeqn{\end{eqnarray}}
\def\no{\nonumber}
\def\nn{\no\\}
\def\eqn#1{(\ref{#1})}
\def\ba{\begin{array}{c}}
\def\ea{\end{array}}
\def\bat{\begin{array}{cc}}
\def\bi{\begin{itemize}}
\def\ei{\end{itemize}}
\def\cA{{\cal A}}
\def\cL{{\cal L}}

\def\cO{{\cal O}}

\begin{document}

\begin{frontmatter}

\title{Nonleptonic $K\to 2\,\pi$ decay dynamics $^*$}

\cortext[cor0]{Talk given at 23th International Conference in Quantum Chromodynamics (QCD 20),  27 - 30 October 2020, Montpellier - FR}

\author{Hector Gisbert}
\ead{hector.gisbert@tu-dortmund.de}
\address{Fakultät für Physik, TU Dortmund, Otto-Hahn-Str.\,4, D-44221 Dortmund, Germany}

\pagestyle{myheadings}
\markright{ }
\begin{abstract}
Using Chiral Perturbation Theory to properly account for the dynamics of nonleptonic $K\to 2\,\pi$ decays, we found the Standard Model prediction for the CP violating ratio $\mbox{Re}\left(\varepsilon'/\varepsilon\right)=\left(14\pm 5\right) \times 10^{-4}$, where isospin breaking effects are included, in perfect agreement with the current experimental world average. Similar results have been reported by a recent release of improved lattice data.
\end{abstract}

\begin{keyword}  

Kaon decays \sep CP violation \sep Standard Model \sep Chiral Perturbation Theory \sep Lattice QCD

\end{keyword}

\end{frontmatter}

\section{Introduction}

The large asymmetry between matter and antimatter in the observable universe requires the existence of additional CP sources beyond the Standard Model (SM). Accordingly, the CP violating ratio $\varepsilon'/\varepsilon$ provides an excellent road towards the discovery of New Physics (NP), since it is linearly proportional to the only existing CP source in the SM. In addition, its study offers a formidable test of flavor-changing neutral-currents (FCNCs) transitions and therefore the Glashow-Iliopoulos-Maiani (GIM) mechanism.

The current experimental world average of $\varepsilon'/\varepsilon$ from NA48~\cite{Batley:2002gn} and KTeV~\cite{AlaviHarati:2002ye,Abouzaid:2010ny},
\begin{equation}\label{eq:exp}
\hspace{1cm}\mathrm{Re}\, (\varepsilon'/\varepsilon)_{\,\mathrm{exp}}\; =\;
(16.6 \pm 2.3) \cdot  10^{-4}\, ,
\end{equation}
exhibits the existence of direct CP violation in decay transitions of $K\to 2\,\pi$, in addition, it is highly sensitive to new sources of CP violation due to its small size.

Since the establishment of its experimental measurement, the SM prediction of $\varepsilon'/\varepsilon$ has been subject of several years of discussions. The first next-to-leading order (NLO) calculations obtained SM values of order $\sim 10^{-4}$~\cite{Buras:1993dy,Buras:1996dq,Bosch:1999wr,Buras:2000qz,Ciuchini:1995cd,Ciuchini:1992tj}, in clear conflict with Eq.~\eqn{eq:exp}. However, it was soon realised by the authors of Refs.~\cite{Pallante:1999qf,Pallante:2000hk,Pallante:2001he}, that an important ingredient had been missed in the previous predictions, the inclusion of final-state interactions (FSI) between the two emitted pions. These corrections were computed within the low-energy effective realization of quantum chromodynamics (QCD), that is Chiral Perturbation theory ($\chi$PT). Once they were included in the calculation, the SM prediction was found consistent with Eq.~\eqn{eq:exp}, albeit with large uncertainties of non-perturbative origin. 

The development of advanced lattice QCD techniques as well as the increasing computer capabilities resulted in a successful calculation of the $\Delta I =3/2$  $K^+\to \pi^+\pi^0$ amplitude \cite{Blum:2011ng,Blum:2012uk,Blum:2015ywa}, and the first statistically-significant signal of the $\Delta I =1/2$ enhancement~\cite{Boyle:2012ys}, both carried out by RBC-UKQCD collaboration and in agreement with previous analytical computations~\cite{Pich:1995qp,Pich:1990mw,Jamin:1994sv,Bertolini:1997ir,Antonelli:1995nv,Antonelli:1995gw,Hambye:1998sma,Bardeen:1986vz,Buras:2014maa,Bijnens:1998ee}. 

Based on these results, the RBC-UKQCD collaboration reported as well the prediction for the CP violating ratio, $\mathrm{Re}\,(\varepsilon'/\varepsilon) = (1.4 \pm 6.9) \cdot 10^{-4}$ \cite{Bai:2015nea,Blum:2015ywa}, with a tension of $\sim 2\,\sigma$ with Eq.~\eqn{eq:exp}. This result revived again the old debate~\cite{Gisbert:2018tuf,Gisbert:2018niu,Gisbert:2019uyx,Buras:2018ozh}, bringing back the past approaches with low $\varepsilon'/\varepsilon$ predictions~\cite{Buras:2015xba,Buras:2016fys,Buras:2015yba}. However, the same lattice study~\cite{Bai:2015nea} reported a $I=0$ phase shift about $\sim 3 \,\sigma$ away from its dispersive prediction, exhibiting that the long-distance re-scattering of the final pions in $K\to 2\,\pi$ was not well-controlled in the simulation. 

Although  the importance of the pion dynamics in the determination of $\varepsilon'/\varepsilon$ was already pointed out in Refs.~\cite{Pallante:1999qf,Pallante:2000hk,Pallante:2001he}, the results required a thorough revision with an update of the different inputs. At the end of 2017, the updated prediction of $\varepsilon'/\varepsilon$~\cite{Gisbert:2017vvj} was found again in good agreement with Eq.~\eqn{eq:exp}.  

Last year, the authors of Ref.~\cite{Cirigliano:2019cpi} performed the update of the last missing ingredient, isospin breaking corrections to $\varepsilon'/\varepsilon$. Including all the sources of uncertainty, the prediction for the CP violating ratio was found in perfect agreement with Eq.~\eqn{eq:exp}. 

The results of Refs.~\cite{Cirigliano:2019cpi} were presented to the high energy physics community in several international conferences~\cite{Cirigliano:2019obu,Cirigliano:2019zjv,Cirigliano:2019ani}, making necessary a revision of the lattice QCD prediction~\cite{Bai:2015nea,Blum:2015ywa} to confirm the findings of Refs.~\cite{Gisbert:2017vvj,Cirigliano:2019cpi}.

After years of efforts to develop the lattice understanding of the pion dynamics, the RBC-UKQCD outcomes~\cite{Abbott:2020hxn} emerged in April 2020, which are outlined in the following~\cite{Blum:2015ywa,Abbott:2020hxn}:
\begin{itemize}
    \item Lattice strong phase shifts in accordance with dispersive predictions.
    \item $K \to 2\,\pi$ isospin amplitudes $A_{0,2}$ in good agreement with their experimental measurements.
    \item $\varepsilon'/\varepsilon$ prediction in agreement with Eq.~\eqn{eq:exp}~,
    \begin{equation}\label{eq:latt}
\hspace{0.5cm}\mathrm{Re}\, (\varepsilon'/\varepsilon)_{\,\mathrm{lattice}}\; =\;
(22 \pm 8) \cdot  10^{-4}\, ,
\end{equation}
    where isospin breaking effects are not included in the central value.
\end{itemize}

These lattice results close the long debate on the SM prediction of $\varepsilon'/\varepsilon$, confirming the observations made 20 years earlier in $\chi$PT~\cite{Pallante:1999qf,Pallante:2000hk,Pallante:2001he,Cirigliano:2003nn,Cirigliano:2003gt}~, and offering good evidence for their updates~\cite{Gisbert:2017vvj,Cirigliano:2019cpi}.\footnote{Using the recent lattice results~\cite{Blum:2015ywa,Abbott:2020hxn}, the authors of Ref.~\cite{Aebischer:2020jto} have also reported a prediction of $\varepsilon'/\varepsilon$ in agreement with Eq.~\eqn{eq:exp}.}

In the following sections, we highlight the essential dynamical features of $K\to 2\,\pi$ decays and briefly summarize the updated $\varepsilon'/\varepsilon$ prediction from Ref.~\cite{Cirigliano:2019cpi}.

\section{Dynamics of $K \to 2\,\pi$ decays}
\label{sec:anatomy}

The physical kaon decay amplitudes can be expressed in terms of isospin components as 
\begin{eqnarray}  
A(K^0 \to \pi^+ \pi^-) &=&  
\cA_{1/2} + {1 \over \sqrt{2}} \left( \cA_{3/2} + \cA_{5/2} \right)  \, ,
\no\\[2pt]
A(K^0 \to \pi^0 \pi^0) &=& 
\cA_{1/2} - \sqrt{2} \left( \cA_{3/2} + \cA_{5/2}  \right) \, ,\no
\\[2pt]
A(K^+ \to \pi^+ \pi^0) &=&
{3 \over 2}  \left( \cA_{3/2} - {2 \over 3} \cA_{5/2} \right)\,,
\label{eq:2pipar}
\end{eqnarray}
with $\cA_{1/2}= A_{0}\, \mathrm{e}^{i \chi_0}$, $\cA_{3/2}= \frac{2}{5}\,A_{2} \,\mathrm{e}^{i \chi_2}+\frac{3}{5}\,A_{2}^+ \,\mathrm{e}^{i \chi_2^+}$, and $\cA_{5/2}= \frac{3}{5}\,A_{2}\, \mathrm{e}^{i \chi_2}-\frac{3}{5}\,A_{2}^+\, \mathrm{e}^{i \chi_2^+}$. $A_0$, $A_2$ and $A_2^+$ are real and positive amplitudes if the CP-conserving limit is assumed. In the isospin limit, $\cA_{5/2}$ is zero, resulting in only two isospin amplitudes $A_{0}$ and $A_2=A_2^+$ corresponding to final states $(\pi\pi)_{I=0,2}$, in addition to its respective phase differences $\chi_0$ and $\chi_2=\chi_2^+$ which represent the S-wave scattering phase shifts. 

Information regarding amplitudes and phase differences can be extracted from the experimental $K\to 2\,\pi$ branching ratios~\cite{Antonelli:2010yf}:
\begin{eqnarray}
A_0 &=& (2.704 \pm 0.001) \cdot 10^{-7} \mbox{ GeV}, \nn
A_2 &=& (1.210 \pm 0.002) \cdot 10^{-8} \mbox{ GeV}, \nn
\chi_0 - \chi_2 &=& (47.5 \pm 0.9)^{\circ},
\label{eq:isoamps}
\end{eqnarray}
which allow us to learn two things about the dynamics of these decays:
\begin{enumerate}
\item Substantial enhancement in the isoscalar amplitude relative to the isotensor one, 
\be
\hspace{1.cm}\omega \,\equiv\, 
\mathrm{Re}\, A_2 / \mathrm{Re}\, A_0\,\approx\, 1/22~.
\ee
\item Large difference between $\chi_0$ and $\chi_2$, which implies by Eq.~\eqn{eq:2pipar} that half of the ratio $\cA_{1/2}/\cA_{3/2}$ is generated from its absorptive contribution:
\be\label{eq:AbsRat}
\mathrm{Abs}\, (\cA_{1/2}/\cA_{3/2})\approx \mathrm{Dis}\, (\cA_{1/2}/\cA_{3/2})~.
\ee
\end{enumerate}
The above statements are general, and represent excellent control tests for $K\to 2\,\pi$ amplitude predictions when a specific theoretical framework is adopted.

In the presence of CP violation, $\mathrm{Im}\, A_{0,2}$ are nonzero and allows us to define the CP violating observable,\footnote{Eq.~\eqn{eq:cp1} is leading order (LO) in CP violation. Corrections of $\mathcal{O}\left((\mathrm{Im}\,A_{0,2})^2\right)$ are very suppressed.}
\begin{equation}
\varepsilon'\,  =\, 
- \frac{i}{\sqrt{2}} \: e^{i ( \chi_2 - \chi_0 )} \:\omega\;\frac{\mathrm{Im}\, A_{0}}{ \mathrm{Re}\, A_{0}}\,
\left[
1 \, - \,\frac{1}{\omega}
\frac{\mathrm{Im}\, A_{2}}{ \mathrm{Im}\, A_{0}} \right] .
\label{eq:cp1}
\end{equation}
Eq.~\eqn{eq:cp1} shows that $\varepsilon'/\varepsilon$ is approximately real ($\chi_2 - \chi_0 - \pi/2\approx 0$), and also that $\varepsilon'$ is suppressed by the ratio $\omega$. In view of the later, one could be worried about an additional suppression from a possible subtle numerical cancellation emerging from 
\begin{equation}\label{eq:xfirst}
    \hspace{2.cm}x\equiv 1 \, - \,\frac{1}{\omega}
\frac{\mathrm{Im}\, A_{2}}{ \mathrm{Im}\, A_{0}}~.
\end{equation}
The study of Eq.~\eqn{eq:xfirst} requires the estimation of $\mathrm{Im}\, A_{0,2}$, which can be done in the limit of a large number of QCD colours where the T-product of two colour singlet currents factorizes
\begin{equation}
  \hspace{1.5cm}\langle Q_{i}\rangle=\langle J\cdot J\rangle\:=\:\langle J \rangle\:\langle J \rangle\:B_i~, 
\label{eq:factorization}
\end{equation}
and the current $\langle J\rangle$ has a well-known representation at LO in the $1/N_C$ expansion. Here, $B_i$ parametrizes the deviation of the true hadronic matrix element $\langle Q_{i}\rangle$ from its large-$N_C$ approximation $\langle J \rangle\:\langle J \rangle$.

Due to the chiral enhancement of $(V-A)\,\times\,(V+A)$ operators, as well as the size of their respective Wilson coefficients, $\mathrm{Im}\, A_{0,2}$ are mainly dominated by $Q_{6,8}$~, respectively. Considering only these two operators, one finds
\beqn
    x\approx 1+\frac{\omega^{-1}}{2\sqrt{2}}\left(\frac{F_K}{F_\pi}-1\right)^{-1}\frac{y_8(\mu)}{y_6(\mu)}\,\widetilde{B}\,
    \approx\, 1-\frac{1}{2}\,\widetilde{B}~,
    \label{eq:xapprox}
\eeqn
with $\widetilde{B}\equiv\frac{B_8^{(3/2)}}{B_6^{(1/2)}}$. $F_K$ and $F_\pi$ are the kaon and pion decay constants, respectively. The CP-odd components of the Wilson coefficients at the short-distance scale $\mu$ are denoted by $y_i(\mu)$~. Since the anomalous dimensions of $y_6$ and $y_8$ are the same at $N_C\to \infty$, the ratio $y_8(\mu)/y_6(\mu)$ is independent of $\mu$ in that limit, showing how the product of two colour-singlet quark currents factorizes, as previously seen in Eq.~\eqn{eq:factorization}.

In the large-$N_C$ limit, $x\approx 1/2$ and therefore $\mathcal{O}(1/N_C)$ corrections are numerically relevant in Eq.~\eqn{eq:xapprox} . Performing an expansion of $B_6^{(1/2)}$ and $B_8^{(3/2)}$ in powers of $1/N_C$, that is $B_i= 1+\frac{1}{N_C}b_i+\mathcal{O}\left(1/N_C^2\right)$~, where $b_i$ are $\mathcal{O}(1)$~, Eq.~\eqn{eq:xapprox} becomes
\beqn
    \hspace{2.cm}x\approx\frac{1}{2}-\frac{1}{3}(b_8-b_6)~,
    \label{eq:xapprox1}
\eeqn
where $N_C=3$ has been employed. Eq.~\eqn{eq:xapprox1} shows explicitly the possibility of a strong cancellation between the two isospin contributions, reaching is maximal efficiency for $b_8-b_6\approx \frac{3}{2}$, however this cancellation depends on the predicted values of $b_6$ and $b_8$ which are $\mathcal{O}(1)$ as already mentioned.

Although isospin breaking effects are suppressed in almost all the phenomenological observables, $\varepsilon'$ represents an exception. In the second term of Eq.~\eqn{eq:cp1}, we observe that a tiny isospin violating correction in $\mathrm{Im}\, A_{2}$ can break its own suppression through the large enhancement of $\omega^{-1}\approx 22$ consequence of the $\Delta I=1/2$ rule. These effects can be naively accounted in Eq.~\eqn{eq:xapprox1}, inserting an additional parameter $\Omega_{\mathrm{naive}}$~, 
\beqn
    \hspace{1.cm}x\approx\frac{1}{2}-\frac{1}{3}\,(b_8-b_6)+\Omega_{\mathrm{naive}}~.\label{eq:naive}
\eeqn
As an example, we observe that LO isospin breaking corrections originated from $\pi^0$-$\eta$ mixing can be estimated as $\Omega_{\mathrm{naive}}\sim \mathcal{O}\left(\omega^{-1}\,\varepsilon^{(2)}\right)\sim\mathcal{O}(10^{-1})$~, where $\varepsilon^{(2)}\sim \mathcal{O}(10^{-2})$ is the tree-level mixing angle. Therefore, isospin corrections in $x$ are relevant and even more if Eq.~\eqn{eq:xapprox1} results to have a strong cancellation.  

For a proper control of the $\varepsilon'/\varepsilon$ prediction, it is useful to rewrite Eq.~\eqn{eq:cp1} to first order in isospin breaking corrections~\cite{Cirigliano:2003nn,Cirigliano:2003gt}
\be\label{eq:epsp_simp}
\mbox{}\hskip -.6cm
\mathrm{Re}\Bigl(\frac{\varepsilon'}{\varepsilon}\Bigr) \, = \, - \frac{\omega_+}{\sqrt{2}\, |\varepsilon|}  \, \left[
\frac{\mathrm{Im}\, A_{0}^{(0)} }{ \mathrm{Re}\, A_{0}^{(0)} }\,
\left( 1 - \Omega_{\rm eff} \right) - \frac{\mathrm{Im}\, A_{2}^{\rm emp}}{ \mathrm{Re}
  A_{2}^{(0)} } \right]  . \;\;
\ee
The superscript $(0)$ denotes the isospin limit, $\mathrm{Im}\, A_2^{\rm emp}$ contains the $I=2$ contribution from the electromagnetic penguin operator, and $\omega_+ \equiv \mathrm{Re}\, A_{2}^{+}/\mathrm{Re}\, A_{0}$~. The parameter $\Omega_{\rm eff}$ encodes isospin-breaking corrections.

To minimize the theoretical uncertainty in our prediction, the CP-conserving amplitudes $\mathrm{Re}\, A_{0,2}^{(0)}$~, and consequently $\omega_+$, are set to their experimental values, provided in Eq.~(\ref{eq:isoamps}). Only $\mathrm{Im}\, A_{0}^{(0)}$, $\mathrm{Im}\, A_{2}^{(0)}$ and $\Omega_{\rm eff}$ (naively related with $b_6$~, $b_8$ and $\Omega_{\mathrm{naive}}$, respectively) require an analytical computation within a theoretical framework. The later has to account for the dynamics of $K\to 2\,\pi$ decays, raised at the beginning of this section, as well as good theoretical control of the different contributions, to disentangle the possibility of cancellation between different contributions as seen in Eq.~\eqn{eq:naive}. The theoretical framework adopted for our prediction is presented in the following section.

\section{Theoretical framework}

The physical origin of $\varepsilon'/\varepsilon$ is at the electroweak scale where all the flavor-changing processes are defined in terms of quarks and gauge bosons. The gluonic corrections to the $K\to2\,\pi$ amplitudes are amplified with large logarithms due to the existence of very different mass scales ($M_\pi < M_K
\ll M_W$), which can be summed up using the Operator Product Expansion (OPE) and the Renormalization Group Equations (RGEs), all the way down to $\mu<m_c$ scales. Finally, in the three-flavor theory, one gets the following effective Lagrangian~\cite{Buchalla:1995vs} 
\be\label{eq:Leff}
\cL_{\mathrm{eff}}^{\Delta S=1}\, =\, - \frac{G_F}{\sqrt{2}}\,
 V_{ud}^{\phantom{*}}V^*_{us}\,  \sum_{i=1}^{10}
 C_i(\mu) \, Q_i (\mu)\, ,
\ee
which is a sum of local operators weighted by $C_i(\mu)$ short-distance
coefficients which depend on the parameters of the heavy masses ($\mu>M$) and the Cabibbo-Kobayashi-Maskawa (CKM) matrix.
At NLO, the Wilson coefficients $C_i(\mu)$ are known~\cite{Buras:1991jm,Buras:1992tc,Buras:1992zv,Ciuchini:1993vr}. This includes both $\cO(\alpha_s^n t^n)$ and $\cO(\alpha_s^{n+1} t^n)$ corrections, where $t\equiv\log{(M_1/M_2)}$ applies to any ratio logarithm with $M_1,M_2\geq\mu$~. Some next-to-next-to-leading-order (NNLO) corrections are already established \cite{Buras:1999st,Gorbahn:2004my} and an attempt for the full short-distance calculation at the NNLO is underway~\cite{Cerda-Sevilla:2016yzo}.  

We can use symmetry considerations below the resonance region, where perturbation theory no longer works, to describe another effective field theory in terms of the QCD Goldstone bosons ($\pi$, $K$, $\eta$). The pseudoscalar octet dynamics are defined by $\chi$PT through a perturbative expansion in powers of momenta and quark masses over the chiral symmetry breaking scale $\Lambda_\chi\sim 1$~GeV. Then, we can build the effective bosonic Lagrangian with the same SU$(3)_L\otimes \:$SU$(3)_R$ transformation properties as Eq.~\eqn{eq:Leff}, leading to $\cL^{\Delta S=1}=\sum_{n=2}^{n_{\mathrm{cut}}}\cL^{\Delta S=1}_n$, where the cut in $n$ depends on how accurate we want to be in our predictions. At LO, only three terms are allowed by symmetries
\begin{eqnarray}\label{eq:LchiPT}
\cL_{2}^{\Delta S=1}\:=\:G_8\:\cL_{8}\:+\:G_{27}\:\cL_{27}\:+\:G_8\:g_{\mathrm{ewk}}\:\cL_{\mathrm{ewk}}\, .
\end{eqnarray}
The chiral low-energy constants (LECs), {\it i.e.} $G_8$, $G_{27}$ and $G_8\, g_{\mathrm{ewk}}$, can not be determined by symmetry principles, however Eq.~\eqn{eq:factorization} allows us to perform a matching between Eqs.~(\ref{eq:Leff}) and (\ref{eq:LchiPT}) in the large-$N_C$ limit, which gives rise to chiral LECs in terms of Wilson coefficients, and so the short-distance dynamics. Since the large-$N_C$ limit is only applied in the matching, missing $1/N_C$ corrections are not enhanced by large logarithms, {\it i.e.} the naive estimation of these contributions is $b_{6,8}^{\mathrm{matching}}\sim\mathrm{log}(\mu/M_\rho)\sim\mathcal{O}(10^{-1})$, one order of magnitude smaller than its natural size, and then numerically irrelevant to play a game in the cancellation of Eq.~\eqn{eq:naive}. In contrast, the decay amplitudes get large logarithmic corrections from pion loops which are a direct consequence of the large phase shift difference between $\chi_0$ and $\chi_2$. The naive estimation of these contributions implies $b_{6,8}\sim \mathrm{log}(\mu/M_\pi)\sim\mathcal{O}(1)$~, and then they are crucial in Eq.~\eqn{eq:naive}. These corrections can be rigorously calculated using the usual $\chi$PT methods~\cite{Pallante:1999qf,Pallante:2000hk,Pallante:2001he}, and they turn out to be positive for $\left. A_{0}\right|_{Q_6}$, while negative for $\left. A_{2}\right|_{Q_8}$, or in other words $b_8<0$ and $b_6>0$. Therefore, the numerical cancellation between the $I=0$ and $I=2$ terms in Eq.~\eqn{eq:naive} is completely destroyed by the chiral loop corrections, leading to a sizeable enhancement of the SM prediction for $\varepsilon'/\varepsilon$, presented in the next section.

\section{\boldmath SM prediction of $\varepsilon'/\varepsilon$}

Taking into account all computed corrections in Eq.~\eqn{eq:epsp_simp}, which we refer for details to~\cite{Gisbert:2017vvj,Cirigliano:2019cpi}, our SM prediction for $\varepsilon'/\varepsilon$ is
\beqn\label{eq:finalRes}
\hspace{-0.7cm}\mbox{Re}\left(\varepsilon'/\varepsilon\right)&\!\!\! =&\!\!\!\left(13.8\pm 1.3_{\,\gamma_5}\pm 2.5_{\,\mathrm{LECs}}\pm 3.5_{1/N_C}\right)\cdot 10^{-4}
\no\\[5pt] &\!\!\! =&\!\!\! 
\left(14\pm 5\right) \cdot 10^{-4}\, .
\eeqn
where isospin breaking corrections have been included, and result in $\Omega_{\rm eff}=0.11\pm 0.09$~\cite{Cirigliano:2019cpi}. The first error represents the uncertainty from NNLO corrections to the Wilson coefficients accounted through the choice of the scheme for $\gamma_5$.  The second error comes from the input values of the strong LECs (mainly $L_{5,7,8}$).  The last error parametrizes our ignorance about $1/N_C$-suppressed contributions in the matching  region  which have been estimated conservatively through the variation of $\mu$ and $\nu_\chi$ in the intervals $[0.9,1.2]$~GeV and $[0.6,1]$~GeV, respectively. 

Our SM prediction for $\varepsilon'/\varepsilon$ is in perfect agreement with the measured experimental value. We have shown the important role of FSI in $K^0\to\pi\pi$. When $\pi\pi$ re-scattering corrections are taken into account, the numerical cancellation between the $Q_6$ and $Q_8$ terms in Eq.~\eqn{eq:naive} is no longer operative because of the positive enhancement of the $Q_6$ amplitude and the negative suppression of the $Q_8$ contribution. 

Although our theoretical uncertainty is still large, improvements can be achieved in the next years:
\begin{itemize}
\item NNLO computation of the Wilson coefficients is close to being finished~\cite{Cerda-Sevilla:2016yzo}. 
\item The $\cO(e^2p^0)$ coupling $G_8\, g_{\mathrm{ewk}}$ can be expressed as a dispersive integral over the hadronic vector and axial-vector spectral functions. The $\tau$ decay data can be used to perform a direct determination of this LEC~\cite{Sanchez:2018atj,AntonioS}.

\item  Estimation of higher-order chiral logarithmic corrections could be feasible either through explicit two-loop calculations or with dispersive techniques~\cite{Pallante:1999qf,Pallante:2000hk,Pallante:2001he,Buchler:2001nm,Buchler:2005xn}.

\item A matching calculation of the weak LECs at NLO in $1/N_C$ remains a very challenging task. A fresh view to previous attempts \cite{Bertolini:1995tp,Bertolini:1997nf,Bertolini:1998vd,Bijnens:2000im,Hambye:2003cy,
Pich:1995qp,Pich:1990mw,Jamin:1994sv,Bertolini:1997ir,Antonelli:1995nv,Antonelli:1995gw,Hambye:1998sma,Bardeen:1986vz,Buras:2014maa,Bijnens:1998ee} could suggest new ways to face this unsolved problem.

\end{itemize}

\section*{Acknowledgements}

I would like to thank Vincenzo Cirigliano, Toni Pich, and Antonio Rodríguez for enjoyable collaborations. This work is supported by the \textit{Bundesministerium für Bildung und Forschung – BMBF}.

\end{document}